1.  A global engineering company accepts a consulting engagement for a firm price of $8,000 in March, does the work in April, bil

2.

On March 31, 2021, Coronado Rental Agency Incs trial balance included the following selected unadjusted account balances. ThDate Debit Credit Account Titles and Explanation | Mar. 31 31 31 31 31 31
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Expenses should be booked in the month incurred.
So $ 4,500 will be booked in the month of April.
Revenue should be booked in the month service provided.
So $ 8,000 will be booked in the month of April.
So, Expenses of $ 4,500 should be deducted from the revenues in April.
Coronado
Working for Adjusting Entry Amount $ Note
Value of equipment 21,600.00 A
Life (Years)             4.00 B
Annual Depreciation Expense     5,400.00 C=A/B
Quarter Depreciation Expense     1,350.00 D=C/12*3
Deferred revenue     9,990.00 E
Portion still unearned 1/3 F
Portion earned 2/3 G=1-F
Portion unearned     6,660.00 H=E*G
Loan payable 18,600.00 I
Interest rate 6% J
Annual Interest amount     1,116.00 K=I*J
Monthly Interest          93.00 L=K/12
Supplies Expense
Supplies - Opening Balance     2,800.00
Less: Supplies - Closing Balance        760.00
Supplies Expense     2,040.00 M
Prepaid Insurance 15,600.00 N
For the period (months)          12.00 O
Monthly Insurance     1,300.00 P=N/O
Insurance from Jan to Mar (3 months)     3,900.00 Q=P*3
Adjusting Entries
Account Debit $ Credit $ Note
Depreciation Expense- Equipment     1,350.00 See D
Accumulated Depreciation- Equipment     1,350.00
Deferred revenue     6,660.00 See H
Rent Revenue     6,660.00
Interest Expense          93.00 See L
Interest Payable          93.00
Supplies Expense     2,040.00 See M
Supplies     2,040.00
Insurance Expense     3,900.00 See Q
Prepaid Insurance     3,900.00
Tax Expense     3,330.00
Tax Payable     3,330.00